OpenStudy (anonymous):

for f(x)=x+4 and g(x)=1/x^2 find (g o f)(4)

OpenStudy (luigi0210):

It's this right?: \[\LARGE f(x)=x+4\] \[\LARGE g(x)=\frac{1}{x^2}\]

OpenStudy (anonymous):

Yes, sorry I did not put it in correctly.

OpenStudy (ranga):

(g o f)(4) = g(f(4)) Find f(4) Then find g(f(4))

OpenStudy (ranga):

f(x) = x + 4 f(4) = ?

OpenStudy (luigi0210):

Ranga's way is easier, but just so you you could do it this way: \[\LARGE g(f(x))=\frac{1}{(x+4)^2}\]

OpenStudy (luigi0210):

After you simplify that, just plug in 4

OpenStudy (anonymous):

Alright thank you!

OpenStudy (anonymous):